# Principal stratification with $U$-statistics under principal ignorability

### v2

@@ 4.1 The efficient influence function for the PGCE
We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al.,, 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis,, 2006; Mao,, 2017). Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators. Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al.,, 2018; Rotnitzky et al.,, 2021). We will explore these properties by first formalizing the EIF for the PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
The EIFs in Theorem 4.1 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey, (2022) and Hines et al., (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ is then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
@@
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression, since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 4.1, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4.1.
@@
with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$. The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$. This coincides with the EIF developed in Mao, (2017) using the projection approach (Tsiatis,, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
@@ 4.2 Multiply robust estimators
The EIFs given in Theorem 4.1 imply the semiparametric efficiency lower bound, and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 4.1, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
@@
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 4.2.
Theorem 4.2 (Multiply robustness and efficiency under parametric working models).
@@
Under Theorem 4.2, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified, and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao, (2017) and Theorem 3 of Jiang et al., (2022). As discussed in Remark 3.2, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies, and is often built upon the orthogonal moment functions (Chernozhukov et al.,, 2018, 2022) or the EIFs. Because of the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al.,, 2022), perturbations in the first-step estimates typically have no significant influence on local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al.,, 2021). Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification. However, machine learning first-step estimates ($\widehat{\pi}(\cdot)$, $\widehat{p}_{z}(\cdot)$, and $\widehat{\mu}_{z_{1}d_{1}z_{2}d_{2}}(\cdot,\cdot)$) have increased complexity and could no longer satisfy the Donsker class condition (Chernozhukov et al.,, 2018). To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al.,, 2018). Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy,, 2023).
Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur, (2023). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $(i,j)$ with $1\leq i<j\leq N$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al., (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
Example 4.3 (Sample splitting for $U$-statistics).
@@
Remark 4.2.
Estimators given in (15) are the type II estimator proposed in Chernozhukov et al., (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
@@
Theorem 4.4 (Efficiency under nonparametric working models).
@@
    there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 4.1.
Theorem 4.4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al., (2022, Lemma 15) and Escanciano and Terschuur, (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 4.4 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Conditions (c) and (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al.,, 2018). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.
@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 3.4, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@ 5.1 The probability index estimands with an intermediate variable
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators, because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 3.2. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified, while the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions). Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
@@
5.2 The win-ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2.2 and 3.3, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:

### v3

@@ 4.1 The efficient influence function for the PGCE
We first develop more efficient estimators with root-$n$ convergence rate under feasible conditions for the PGCEs by obtaining their corresponding EIFs (Bickel et al., 1993). The EIF is a nonparametric function of observed data that characterizes the target estimand, and its form is instrumental in developing optimal estimators. Several key properties of the EIF suggest its promise for constructing optimal estimators. First, the EIF often integrates more than one nuisance model and facilitates estimators that are robust against misspecification of some nuisance models (Tsiatis, 2006; Mao, 2017). Second, the asymptotic variance lower bound is represented by the variance of the EIF, providing a useful benchmark to define efficient estimators. Third, the EIF provides a device for integrating advanced machine learners for the nuisance parameters, which can lead to feasible estimators that achieve the asymptotic variance lower bound (Chernozhukov et al., 2018; Rotnitzky et al., 2021). We will explore these properties by first formalizing the EIF for the PGCE. Because the PGCEs are of a ratio structure, we define $\tau_{h}^{s}=\tau_{h,N}^{s}/\tau_{D}^{s}$ for $s\in\{10,00,11\}$, where, for example, $\tau_{h,N}^{10}=\mathbb{E}\{e_{10}(X_{1})e_{10}(X_{2})\allowbreak\mu_{1100}(X_{1},X_{2})\}$ and $\tau_{D}^{10}=(p_{1}-p_{0})^{2}$. This suggests that the EIF for $\tau_{h}^{s}$, $\varphi_{h}^{s}(\mathcal{O})$, is of the form:
@@
The EIFs in Theorem 2 are derived using the method of Gâteaux derivatives as formalized in Ichimura and Newey (2022) and Hines et al. (2022) with respect to the distribution of observed data, $\mathcal{O}$. The derivation for $\varphi_{D}^{s}(\mathcal{O})$ then follows the chain rule. For $\varphi_{h,N}^{s}(\mathcal{O})$, we re-express the PGCE in terms of the observed data as
@@
where $\mathfrak{h}_{1}^{s}(y)=\mathbb{E}\{h(y,Y^{\prime})|Z^{\prime}=0,S=S^{\prime}=s\}$ and $\mathfrak{h}_{0}^{s}(y)=\mathbb{E}\{h(Y^{\prime},y)|Z^{\prime}=1,S=S^{\prime}=s\}$, with the conditional expectations taken with respect to $Y^{\prime}$. Importantly, the EIF for the PGCE remains unchanged after this re-expression since the definition of the estimand, the average pairwise contrast between treatment and control conditions, and the data-generating process remain unchanged. Based on Theorem 2, the asymptotic variances of estimators for the PGCEs are lower bounded by the semiparametric efficiency bounds, $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})^{2}\}$, as $\mathbb{E}\{\varphi_{h}^{s}(\mathcal{O})\}=0$ for $s\in\{10,00,11\}$.
Remark 4
@@
with $\nu_{z}(Y)=\int h(Y,y)\mathrm{d}F_{z}(y|x)\mathrm{d}F_{X}(x)$, $m_{1}(X)=\int\mu_{10}(X,x)\mathrm{d}F_{X}(x)$ and $m_{0}(X)=\int\mu_{10}(x,X)\allowbreak\mathrm{d}F_{X}(x)$. The function $\mu_{10}(X_{1},X_{2})=\mathbb{E}\{h(Y_{1},Y_{2})|Z_{1}=1,Z_{2}=0,X_{1},X_{2}\}$ is the pairwise outcome mean within the observed stratum $(Z_{1}=1,Z_{2}=0)$. This coincides with the EIF developed in Mao (2017) using the projection approach (Tsiatis, 2006), and intuitively, $\varphi_{h}(\mathcal{O})$ can be viewed as a special case of $\varphi_{h,N}^{s}(\mathcal{O})$ in the absence of the intermediate variable.
@@ 4.2 Multiply robust estimators
The EIFs given in Theorem 2 imply the semiparametric efficiency lower bound and pave the way for formulating estimators for the PGCEs. However, each EIF involves unknown marginal functions such as $m_{z}^{s}(\cdot)$ and $\nu_{1,00}^{10}(\cdot)$ that are defined as integrals of the pairwise mean function and requires care for constructing the final estimator. We consider the following $U$-statistics in constructing the final estimator, which could be shown to have the same influence functions as those given in Theorem 2, thereby achieving semiparametric efficiency with correctly specified nuisance functions. To this end, we first notice that each PGCE can be rewritten as the following:
@@
which are essentially $U$-statistics but in the form of ratios. Their asymptotic properties are given in Theorem 3.
Theorem 3 (Multiply robustness and efficiency under parametric working models)
@@
Under Theorem 3, $\widehat{\tau}_{h}^{s}$ is consistent if any two of the three parametric nuisance functions are correctly specified and is semiparametrically efficient if all three nuisance models are correctly specified. The triple robustness property of $\widehat{\tau}_{h}^{s}$ is obtained by separately investigating its numerator and the denominator. The denominator is consistent for $e_{s}^{2}$ under $\mathcal{M}_{\pi}\cup\mathcal{M}_{p}\subset\mathcal{M}_{\pi+p+\mu}$. The consistency of the numerator is more involved and requires calculating the asymptotic bias of $\mathbb{U}_{n}g_{h}^{s}(\widehat{\theta})-\tau_{h,N}^{s}$. The specific derivations are provided in the Online Supplement Section S4.1 due to the complexity of the expressions. The proof of semiparametric efficiency of $\widehat{\tau}_{h}^{s}$ is built upon Lemma S1 of Mao (2017) and Theorem 3 of Jiang et al. (2022). As discussed in Remark 3, the asymptotic behavior of $\widehat{\tau}_{h}^{s}$ can be investigated by studying that of the denominator and the numerator after applying the Taylor’s expansion on the ratio estimator. Details of the proof are provided in Section S4.2 of the Online Supplement. Finally, The variances of the triply robust estimators given in (14) can be estimated via nonparametric bootstrap.
@@ 4.3 Debiased machine learning estimators
Debiased machine learning is a burgeoning method for efficient estimation of the causal effects in experimental and observation studies and is often built upon the orthogonal moment functions (Chernozhukov et al., 2018, 2022) or the EIFs. Because the Neyman orthogonality is satisfied by the EIF (Chernozhukov et al., 2022), perturbations in the first-step estimates typically have no significant influence on the local properties of the desired estimator for the causal estimand. In addition, the mixed bias property of the EIF representation allows one to obtain $\sqrt{n}$-consistent and asymptotically normal estimators when one can estimate a subset of nuisance functions at sufficiently fast rates (Rotnitzky et al., 2021). Compared to parametric working models, leveraging machine learning working models for the nuisance functions is more advantageous for its robustness against potential misspecification. To control the remainder empirical process terms of the debiased machine learning estimators to be asymptotically negligible, a sample-splitting procedure—cross-fitting is adopted, where the entire sample is divided into separate parts for the estimation of the nuisance functions and the estimation of the causal estimand respectively (Chernozhukov et al., 2018). Mild regulatory conditions, mostly the consistency of machine learning first-step estimates, are required for the $\sqrt{n}$-consistency and asymptotic normality of the resulting debiased machine learning estimator (Kennedy, 2023).
Because of the unique structure of the PGCEs, we adopt the sample-splitting scheme for $U$-statistics provided in Escanciano and Terschuur (2023). Specifically, the sample index set $\mathcal{N}=\{1,\ldots,n\}$ is first divided into $K$ subsets of similar sizes, forming a partition =$\mathcal{I}=\{I_{1},\ldots,I_{K}\}$. Then, upon $\mathcal{I}$, a partition, $\mathcal{P}$, can be formed for all pairs $\{(i,j),1\leq i<j\leq N\}$, which has $L=K(K+1)/2$ elements, i.e., $\mathcal{P}=\{P_{1},\ldots,P_{L}\}$. Chernozhukov et al. (2018) suggested that $K$ be set to moderate values such as 4 or 5, rendering sizes of elements in $\mathcal{I}$ are of the same order as $\mathcal{N}$. The following example is used to provide a demonstration of $\mathcal{I}$ and $\mathcal{P}$.
Example 6 (Sample splitting for $U$-statistics)
@@
Remark 5
Estimators given in (15) are the type II estimator proposed in Chernozhukov et al. (2018) for estimands without an intermediate variable. Alternatively, one could first obtain separate estimates
@@
Theorem 4 (Efficiency under nonparametric working models)
@@
    $g_{h,1}^{s}(\mathcal{O};\theta)\equiv\mathbb{E}g_{h}^{s}(\mathcal{O},o;\theta)$ and $\psi_{D_{z}}(\mathcal{O};\vartheta)$ are $\mathbb{P}$-Donsker with probability approaching 1, and there exists $C>0$ such that $|\mathbb{E}g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s*}|\leq C\|\theta-\theta^{*}\|^{2}$ and $|\mathbb{E}\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}^{*}|\leq C\|\vartheta-\vartheta^{*}\|^{2}$ for $\theta,\vartheta\in\Theta^{*}$ when $\|\theta-\theta^{*}\|$ and $\|\vartheta-\vartheta^{*}\|$ are small,
then $\widehat{\tau}_{h}^{s}$ in (15) is $\sqrt{n}$-consistent, asymptotically normal, and has influence function $\varphi_{h}^{s}$, thereby achieving the semiparametric efficiency lower bound characterized by the EIF in Theorem 2.
Theorem 4 establishes the consistency and asymptotic normality of the debiased machine learning PGCE estimators, building on the results developed for orthogonal moment functions given in Chernozhukov et al. (2022, Lemma 15) and Escanciano and Terschuur (2023, Lemma 4). To invoke previously established results, it is required that both score functions, $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)-\tau_{h,N}^{s}$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)-p_{z}$, satisfy the Neyman orthogonality, which can be verified by showing $\mathbb{E}\{\partial g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)/\partial\theta\}=0$ and noting that $\varphi_{p_{z}}(\mathcal{O})\equiv\psi_{D_{z}}(\mathcal{O};\vartheta)-p_{z}$ is the EIF for $p_{z}$. Regulatory conditions in Theorem 4 are mostly mild. Condition (a) is a standard boundedness requirement on functions $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta^{*})$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta^{*})$. Condition (b) assumes that machine learning first-step estimators are consistent as well as the $\mathcal{L}_{2}$-moment convergence of $g_{h}^{s}(\mathcal{O}_{i},\mathcal{O}_{j};\theta)$ and $\psi_{D_{z}}(\mathcal{O}_{i};\vartheta)$ are attainable, which are generally achievable for commonly used machine learners. Condition (c) and part of condition (d) together constitute the small bias assumption for $\widehat{\theta}_{l}$ and $\widehat{\vartheta}_{k}$; Condition (c), in particular, specifies the required convergence rate for first-step estimates at $n^{-1/4}$, which could be satisfied by most existing data-adaptive machine learning methods (Chernozhukov et al., 2018). The $\mathbb{P}$-Donsker class assumption in condition (d) regulates the complexity of the function classes for both the numerator and denominator, which has been commonly adopted in debiased machine learning literature (Vansteelandt and Dukes, 2022; Wang et al., 2024). Finally, for statistical inference, we estimate the variances of the debiased machine learning estimators given in (15) via nonparametric bootstrap.
@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 1, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@ 5.1 The probability index estimands with an intermediate variable
to simulate scenarios where the nuisance functions are misspecified. Specifically, we fit the model using $X_{i}$ in scenarios where the respective nuisance function is correctly specified, and $\widetilde{X}_{i}$ otherwise. Also, it is worth noting that, in practice, model misspecification pertains to the flexible debiased machine learning estimators because their estimation accuracy could be affected by the quality of the input covariate feature matrix.
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
Figure 1 presents the results of the five estimators under comparison in scenarios where, at most, one of the three nuisance functions is misspecified. Plug-in moment estimators are consistent whenever the relevant nuisance functions are correctly specified. In contrast, the triply robust and debiased machine learning estimators are consistent across all four scenarios (correct specification of all nuisance functions or subsets of nuisance functions). Results of each considered estimator in scenarios where at least two of the three nuisance functions are misspecified are presented in Web Figure 1 in Section S6 of the Online Supplement. As expected, the plug-in moment estimators are generally biased in those scenarios. In comparison to the plug-in moment estimators, although biased, the triply robust and debiased machine learning estimators overall exhibit improved estimation accuracy. Finally, compared to the triply robust estimator, the debiased machine learning estimator leads to a slight improvement and is generally less biased.
@@
5.2 The win ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model: